### v2

# Adaptive Manipulation using Behavior Trees

Abstract
Many manipulation tasks pose a challenge since they depend on non-visual environmental information that can only be determined after sustained physical interaction has already begun. This is particularly relevant for effort-sensitive, dynamics-dependent tasks such as tightening a valve. To perform these tasks safely and reliably, robots must be able to quickly adapt in response to unexpected changes during task execution. Humans can intuitively respond and adapt their manipulation strategy to suit such problems, but representing and implementing such behaviors for robots remains an open question. We present the adaptive behavior tree, which enables a robot to quickly adapt to both visual and non-visual observations during task execution, preempting task failure or switching to a different strategy based on data from previous attempts. We test our approach on a number of tasks commonly found in industrial settings. Our results demonstrate safety, robustness (100% success rate for all but one experiment) and efficiency in task completion (eg, an overall task speedup of 46% on average for valve tightening), and would reduce dependency on human supervision and intervention.
I Introduction
Service robots can revolutionize how entire service-oriented domains work, with potential applications including autonomous maintenance in – often remote – industrial facilities [1, 2], such as oil rigs [3], or care robots in domestic environments [4, 5, 6]. To complete tasks effectively, robots must use observations from the environment to construct and execute a suitable plan; a number of frameworks have been developed for task and motion planning (TAMP) for manipulation [7, 8, 9, 10].
This work is motivated by effort-sensitive, dynamics-dependent manipulation tasks for articulated devices commonly found in industrial settings, for example twisting a valve until tightened to a specified torque or pushing an e-stop button until pressed. Robots generally require highly controlled environments with known task parameters for reliable planning and execution of manipulation tasks; however, the tasks we are interested in pose a significant challenge since they depend on information that can only be determined after sustained physical interaction has already begun. For example, a robot cannot determine the initial tightness of a valve or stiffness of a button from visual data alone, and can only gather this information by beginning to interact with it. Even then, in the absence of an accurate model for the dynamic profile for the device, the robot cannot know for how long it must twist the valve, or how hard it must push the button, to complete the task. Even if provided a model, stiction from corrosion or unclear servicing may lead to uncertainty and deviation from standard assumptions. To solve these kinds of tasks, we argue that the robot must be able to instead decide on an initial plan with which to start the task and adapt to observations it makes during task execution. Since the tasks we consider may require the robot to apply a significant amount of effort to manipulate the device, adaptability is not only critical for task completion but also for safe robot operation; failure to respond to a sudden unexpected spike in effort can cause catastrophic failure and even damage to the robot or environment.
Existing TAMP frameworks generate plans based on the current task environment, enabling robots to adapt to uncertain and changing environments; however, this adaptability is restricted to visually observable changes, such as for real-time collision avoidance. We require a solution that can handle unexpected deviations that are not visually observable.
We also consider the utility of a robot having access to different manipulation strategies for the same task. A strategy is the specific approach that a robot may take to carry out a task. For manipulation tasks, different strategies may involve grasping the object differently, or even using a tool to assist manipulation. Different strategies can have different trade-offs; for example speed, complexity, effort and power expenditure. For example, quickly grasping and twisting a valve handle head-on may be optimal when beginning to twist an initially loose valve, but fail to provide the torque to finish tightening it, while grasping from the side to apply more wrench may be slower for most of the tightening procedure but can fully tighten the valve. It may even be that completely different strategies are required to successfully complete different task instances. We seek a system to coordinate this strategy selecting and switching behavior.
We tackle both problems by exploiting the reactivity and modularity of behavior trees. The key contribution of our work is the adaptive behavior tree, which enables a robot to handle uncertainty in the environment and quickly adapt to unexpected safety-critical changes during task execution, preempting task failure or switching to a more optimal strategy based on data gathered over previous attempts. It comprises two stages; a selector that uses data recorded over previous attempts to determine which strategy to deploy, and a monitor that checks a set of conditions concurrently to strategy execution, halting the current strategy if the task is complete or a safety condition is violated and re-attempting the task if necessary. The design is highly modular and readily allows for the addition and swapping out of strategy subtrees and condition nodes to suit a new task. We test our approach on various case studies representing complex, effort-intensive manipulation tasks for an industrial setting.
II Related Work
II-A Adaptive Control for Robotics
It may be argued that our problem of manipulating articulated devices with unknown and changing dynamics intrinsics can be solved with an appropriate adaptive controller. Adaptive control has existed in the literature for many years, and generally focuses on online parameter estimation and learning for uncertain dynamic systems. There exist many good surveys on the literature for adaptive control [11, 12]. The approach has been extended to robotic manipulation tasks, for example estimating model parameters for an uncertain end-effector payload [13, 14, 15].
However, we argue that adaptive behavior by parameter tuning is inherently limited in the diversity that it provides for adaptability, and does not readily allow for adaptability to the extent of reactively switching to a completely different manipulation strategy, such as grasping the device in a different way or using a tool, which may be necessary for task completion. Thus, the adaptive behavior we seek must go beyond the control level to the planning level; we opt to instead tackle the problem as a task planning problem. Doing so also allows us to define a more diverse range of conditions for task success and when to halt a strategy, and how to handle when these conditions are met. We will see that our approach can enable a robot to achieve impressive task performance despite only using out-the-box controllers.
II-B Behavior Trees for Robotics
There is much work on task planning and autonomous decision-making for robotics, with the state of the art typically adopting a Finite State Machine (FSM), Behavior Tree (BT) or Partially Observable Markov Decision Process (POMDP) [16, 17]. The appropriate choice depends on the application and intended robot behavior.
A BT is a hierarchical directed tree that controls the order of execution of all the tasks that the robot may perform, each represented as a leaf node. Subtrees representing simpler behaviors can be used to build more complex ones, giving BTs a very composable and modular structure. During normal operation a ‘tick’ signal is propagated between the root of the tree and its leaves, with the active leaf node either returning Success, Failure or Running. This return signal determines which node is executed on the subsequent tick. It can be demonstrated that BTs offer far more modularity than FSMs [17, 18, 19], and BTs also offer superior reactivity by allowing for the concurrent execution of different tree nodes that can interrupt each other if needed. BTs will be our planning framework of choice since the reactivity that they offer is critical for the manipulation tasks we consider, and our approach will also exploit their modularity.
There are many examples demonstrating how BTs can be used to coordinate manipulation tasks [20, 21, 22] and robot control [23, 24, 25]; Iovino et. al. provide a comprehensive survey of BTs for robotics [16]. BTs for robotics can be hand-coded [26, 27, 28], planned analytically [29, 30, 31, 32, 33, 34], or even learned from demonstration [35, 36].
Reactive BTs are particularly relevant to our problem; they use a planning algorithm to automatically create and update a BT on the fly based on observations during task execution, achieving robot behavior that is adaptive to uncertain and changing environments [29, 30]. While this is very promising work, due to the safety-critical nature of the manipulation tasks we aim to tackle, we will opt for an approach that provides a more predictable method of handling failure preemption and recovery.
On the other hand, Wu et. al. [22] propose an automated BT error recovery framework that uses a Bayesian Network (BN) as a decision model for error recovery upon task failure, alongside a human-made Failure Mode and Effects Analysis (FMEA) chart used to inform the network of possible sources of error as well as the corresponding solutions. In contrast, we desire a robot behavior that can preempt failure entirely by stopping and re-attempting the task once a safety condition is violated, potentially using a different strategy.
III Behavior Trees – An Overview
Without loss of generality, we follow the BT notation used by BT.CPP [37]; all BT nodes used in this work fall under one of a set of categories represented in Fig. 1.
A ControlNode controls the sequence of execution of its children; those used in this work are of type Sequence, ReactiveSequence, Fallback or ReactiveFallback.
A Sequence propagates the tick to its children from left to right, returning Success if and only if all their children return Success. When it receives Running, it returns Running and ticks the same child on the next tick, and when it receives Failure, it returns Failure and resets the sequence progress. A ReactiveSequence behaves similarly, but when it receives Running, it returns Running and ticks all children up to the current child on the next tick. Thus, all prior children are still checked for Failure whenever it is ticked; this allows for reactive behavior.
A Fallback (known as a Selector in other works) propagates the tick to its children from left to right until one of them returns Success, at which point the Fallback stops and returns Success. When it receives Running, it returns Running and ticks the same child on the next tick. It returns Failure only if all children return Failure. A ReactiveFallback behaves similarly, but when it receives Running, it returns Running and ticks all children up to the current child on the next tick. Thus, all prior children are still checked for Success whenever it is ticked; this allows for reactive behavior much like in the case of a ReactiveSequence.
An ExecutionNode carries out a command, forming the leaves of the BT, and can be of type ActionNode or ConditionNode. An ActionNode performs a command, returning Success if the command is successfully completed, Failure if failed, or Running if ongoing. A ConditionNode checks a condition, returning Success if true or Failure if false.
A DecoratorNode modifies the return status of a single child node, for example forcing Failure. A Precondition is a special type of DecoratorNode that ticks the child only if a particular condition is met, or else returns either Success or Failure (marked by ‘S’ or ‘F’ respectively).
Figure 1: Categories of BT tree nodes used in this work. From left to right; ControlNode (Sequence, ReactiveSequence, Fallback, ReactiveFallback), ExecutionNode (ActionNode, ConditionNode), DecoratorNode, Precondition, BT.
IV Adaptive BT
Fig. 2 presents the proposed design for an adaptive BT, as well as an example of how it can be integrated within a wider robot system. The adaptive BT is composed of two core stages; the strategy selector first decides which strategy to use for the next attempt, based on current perception data as well as data recorded over past attempts (such as joint efforts over time). The condition monitor then wraps around the subtree for the selected strategy, checking a set of task- and strategy-specific conditions concurrently to strategy execution. The adaptive BT also admits subtrees to coordinate robot setup prior to strategy selection, and for safely exiting or recovering from the current attempt in preparation for the next.
Figure 2: Adaptive BT design, integrated within a wider robot system representing the test platform in this work. The coloured lines represent communication between the BT and the rest of the robot.
The database of the subtrees for all $N\in\mathbb{N}$ known strategies are inserted as modules within a Sequence, but all are preconditioned to only be executed if that specific strategy has been selected. Thus, if strategy $n$ is selected, only that strategy’s subtree will be ticked, and all the other subtrees trivially return Success. If no strategies are considered feasible for the current task instance, SelectStrategy returns Failure and the adaptive BT exits as failed.
The condition checking is such that if the subtree of task-specific conditions returns Success, the task is considered to have been completed successfully, and if it returns Failure, the task is still ongoing. This follows the BT design principle of backward chaining, commanding the robot to keep executing the strategy until the task completion conditions are met. Meanwhile, if the subtree of strategy-specific conditions returns Failure, a constraint has been violated during execution of the current strategy and the robot must halt the current attempt and try again, potentially selecting a different strategy. If it returns Success, no violations have occurred. This logic makes the condition monitor reactive and interpretable. If there are multiple task- or strategy-specific conditions to check, they should be placed in a ReactiveSequence to ensure proper operation of the adaptive BT. Each condition is ticked for every tick of the strategy subtree, allowing the robot to rapidly respond to new observations; thus the inherent reactivity of BTs is exploited. If a condition is not relevant to the current stage of the strategy subtree, we recommend skipping it by wrapping with an appropriate Precondition.
The subtrees for exit and recovery allow the robot to safely exit the current attempt, due to task success or preempting failure respectively. A ForceFailure is wrapped around the recovery subtree to signal that the task is not yet complete and a re-attempt is required.
Very little has been assumed concerning the algorithm behind how the next strategy is selected, or the implementation of any of the subtrees within the strategy selector or condition monitor. The adaptive BT is thus highly modular, and generalizes to a wide range of different problems and can satisfy all manner of tasks, specifications and constraints; all that is required are the subtrees for the task and strategy.
The adaptive BT can easily be added as a subtree within a greater BT structure, and is thus readily applicable to existing BT-based robot task planning systems without major changes. The only required knowledge for adding the proposed design into a larger BT is that the adaptive BT must be repeatedly ticked while Running, and will return Success if the task was completed successfully and Failure otherwise. To avoid getting stuck, we recommend limiting the adaptive BT to a finite number of allowed attempts.
V Adaptive BT for Manipulation
We now apply the proposed adaptive BT to the context of manipulation tasks wherein the robot must manipulate an articulated device; task success occurs when the device is considered sufficiently manipulated. For all manipulation tasks explored in this work, the manipulation procedure consists of identifying the pose of the device, then approaching, grasping, manipulating and finally retracting from the device. Fig. 3 presents the designs used for each subtree of the adaptive BT. We stress that these designs are just examples used to test the adaptive BT, with an emphasis on simplicity.
Figure 3: Subtrees used for all manipulation tasks considered in this work.
Setup. The robot moves to a pose suitable for observing the task instance to determine which strategy to select, and then estimates the pose of the device handle. If the device handle has rotational symmetry and thus has more than one valid pose, the robot later uses the most suitable pose that is within safe device handle pose limits for the selected strategy.
Strategy Subtree. The robot approaches, grasps and manipulates the device with its end-effector. These motions are each defined by a waypointed end-effector pose trajectory that is solved using the robot’s motion planner, with the planning solution then sent to the robot’s controllers. The pose trajectories are different between strategies, enabling the robot to manipulate the device in different ways. While the strategies tested in this work all use the same subtree design, this is not a limitation of the adaptive BT, and in general each strategy can use a completely different design, enabling greater variety between strategies.
Task-Specific Conditions. If currently manipulating the device, the robot checks whether the device is considered sufficiently manipulated. This condition is specified differently depending on the particular task; for example, when tightening a valve, the current joint efforts are used to estimate tightness, while when pressing a button, the joint efforts are used to estimate whether it has been pressed.
Strategy-Specific Conditions. If approaching, grasping or manipulating the device, the robot checks that its end-effector pose is within tolerance of the current pose trajectory. When manipulating the device, the robot checks that its joint efforts do not exceed the safe limits for the strategy.
The robot also checks that the device handle pose is still within safe limits for continuous manipulation using the strategy. This is important since a task like tightening a valve may require multiple revolutions of the handle, yet the robot may not be physically able to do this in one motion due it its own joint limits. Thus, this condition tells the robot when it should stop and re-grasp the device handle, preventing the robot from overextending itself during manipulation.
Recovery/Exit. Both subtrees use the same design; the robot retracts its end-effector from the device (if grasped) and then resets to its starting pose and joint configuration. The retraction is also defined by a waypointed end-effector pose trajectory and solved via motion planning.
VI Implementation
Test Platform. We used an ANYbotics ANYmal C quadruped[38] with mounted Unitree Z1 robotic arm[39], equipped with a wrist-mounted Intel Realsense Depth Camera D345[40] for perception. ROS, the Robot Operating System, was used for the implementation of the test setup as well as communication with hardware [41].
Behavior Tree. BT.CPP 4.0 was used to construct the adaptive BT. The ActionNode types specific to this work take the form of ROS action and service clients, implemented as derived classes of the StatefulActionNode from the BT.CPP core library; this node type supports asynchronous actions and is recommended for this use case.
Strategy Selection. The SelectStrategy node estimates the optimal manipulation strategy. For this work, the optimal strategy $a^{*}$ completes the task in the shortest time while being feasible given the effort data recorded during manipulation. This is equivalent to the following problem:
$a^{*}=\min_{a\in\mathcal{A}_{\mathrm{tsk}}}t_{\mathrm{est}}(a)\quad\text{s.t.}\quad T_{\mathrm{max},i}(a)<T_{\mathrm{lim},i}(a)\;\forall i$ (1)
where strategies $a$ are selected from the set of available task strategies $\mathcal{A}_{\mathrm{tsk}}$, $t_{\mathrm{est}}(a)$ is the estimated time for task completion, $T_{\mathrm{max},i}(a)$ are the largest recorded absolute effort values for each arm joint $i$, recorded from previous manipulation attempts using $a$, and $T_{\mathrm{lim},i}(a)$ are the maximum permitted absolute effort values for each joint $i$ under strategy $a$.
The effort data used for strategy selection is recorded from the arm joint efforts during each manipulation attempt, and stored in a database accessible to SelectStrategy.
Motion Planning. This work uses a whole-body motion planner made using EXOTica, the EXtensible Optimisation Toolset [42]. The planner solves an EXOTica DynamicTimeIndexedShootingProblem, tracking pre-defined waypointed end-effector pose trajectories for the approach, grasp, manipulate and retract stages of the manipulation task for each strategy. The pose trajectories are defined relative to the pose of the device handle; in this way, the planner is generalizable to arbitrary device poses and handle angles (so long as robot joint limits permit the motion). In the case of planning failure, the planner action nodes return Failure, prompting the adaptive BT to retry the task.
Control. This work uses the standard controllers provided with the ANYmal C and Unitree Z1. These controllers are by no means perfect, and in practice there was often non-negligible deviation between the true end-effector pose and that in the motion plan, especially when the arm was outstretched; to ensure safe operation, this deviation is checked by the condition monitor during manipulation. This is a major cause of the re-attempts that would occasionally occur during testing as seen in section VIII. Using a bespoke whole-body controller that compensates for both the motion of the base and arm would improve performance, but this is beyond the scope of this work; instead, we demonstrate how our adaptive BT can be used to overcome these problems by autonomously preempting failure and retrying.
Perception. An Apriltag [43] was used to identify the pose of the device handle in an ‘unrotated’ state, and invariant template matching [44] was used to identify the handle angle (and thus the pose of the rotated handle).
VII Experiments
Needle Valve. The robot twists a Tameson NLS-012 needle valve[45] from unknown starting tightness until tight. Two strategies were provided; the faster, simpler ‘low-torque’ strategy grasps the handle head-on while the slower, more complex ‘high-torque’ strategy grasps it from the side. Fig. 4 depicts these strategies. The experiment is conducted using each strategy on its own, as well as using both strategies such that the BT can adaptively select between them.
To demonstrate the modularity offered by the adaptive BT, we also consider a walk-and-manipulate task where the robot must navigate to, tighten, and navigate away from the needle valve. The adaptive BT is inserted as a subtree within a larger BT that also commands navigation to and from the device (using the standard ANYmal C navigation and locomotion software). Fig. 5a-5c illustrates this experiment.
Figure 4: An annotated timeline of tightening the needle valve using the adaptive BT with both available strategies, with plots of the effort in the wrist joint during each manipulation attempt A, B, C and D. Note the effort spikes during attempts C and D; the first indicates that the low-torque strategy is no longer feasible, while the second indicates that the device is sufficiently tightened. For clarity, only the lowest active effort limit is shown in each plot.
Figure 5: Snapshots of experiments performed. (5a) Navigating to the needle valve. (5b) Tightening the needle valve. (5c) Navigating away from the needle valve. (5d) Pushing the e-stop button until fully pressed. (5e) Twisting the e-stop button until released. (5f) Tightening the globe valve.
E-Stop Button. The robot manipulates an IDEC HW1X-BV411-R emergency stop[46]; it pushes the button until fully pressed and twists it until released. One strategy was provided for each task; Fig. 5d and 5e depict these strategies. A couple thin foam strips were placed on the Unitree Z1’s gripper to allow it to tightly grasp the button for twisting.
Globe Valve. The robot twists a Dikkan DN40 PN16 globe valve[47] from unknown starting tightness until tight. One strategy was provided for this task; Fig. 5f depicts this strategy. The valve’s diaphragm and O-ring were removed to make it loose enough to be twisted by the Unitree Z1, and some thin foam strips were placed on the handle to allow the Z1’s gripper to tightly grasp the handle.
VIII Results and Discussion
An accompanying video for the experiments performed is provided with this work [48].
Each experiment was repeated for 10 trials with the device handle pose and recorded effort data reset for each trial. Strategy effort limits were set to the Unitree Z1’s maximum safe operating efforts scaled down by a small safety factor, except for the low-torque strategy which used a larger safety factor to avoid the base becoming unstable from the arm’s wrench during heavy twisting. Task success was verified by checking the device after the robot had exited the task. Table I records, for each experiment, the cumulative success rate after each attempt, ignoring re-attempts triggered to change strategy and re-grasping when device handle pose limits were exceeded, as well as times taken for task completion. A maximum of 5 attempts was permitted for each trial. Fig. 4 provides an annotated timeline of tightening the needle valve when having access to both associated strategies.
TABLE I: Results for All Manipulation Tasks Using the Adaptive BT
Device Experiment Cumulative Success Rate (%) After Attempt Number Time Taken (s)
1 2 3 4 5 Minimum Mean
Needle Valve Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 135 (Fail) 136 (Fail)
High-Torque Strategy Only 50 70 90 100 Done 279 336
Both Strategies (Adaptive) 80 100 Done Done Done 175 182
Walk-and-Manipulate 60 80 90 100 Done 212 230
E-Stop Button Press 80 100 Done Done Done 28 32
Release 70 80 100 Done Done 45 63
Globe Valve Tighten 70 100 Done Done Done 95 107
Table I shows that the robot achieved a 100% success rate within the permitted number of attempts for all but one experiment when using the adaptive BT, despite having no knowledge of the valves’ initial tightness, or how much they needed to be twisted to be tightened. The robot never suffered catastrophic failure since the condition monitor preempted failure before it occurred. When only having access to the low-torque strategy, the robot aborted tightening the needle valve after the measured joint effort exceeded the strategy effort limit for the arm’s wrist joint. This demonstrates the safety enabled by the adaptive BT, since the robot was able to react and safely exit the task before overexerting itself, and would not try again because the manipulation data suggested that no feasible strategies were available. However, the utility of having access to the low-torque strategy is made clear when comparing the task completion times for using the high-torque strategy only and adaptively using both strategies. Since the low-torque strategy is approximately 5 times faster while twisting, it is optimal for the robot to use it for most of the manipulation sequence, but when the robot experiences a sudden effort spike as the valve begins to tighten, it quickly adapts and uses the high-torque strategy to fully tighten the valve; this resulted in an overall task speedup of 46% on average. We note that this is similar to behaviour displayed by humans, twisting quickly when the valve is loose but changing grip near the end to be able to apply more wrench without hurting oneself. The robot was also able to safely twist the valves for an arbitrary number of rotations without the danger of overextending itself. This was done by periodically re-grasping the valves using the condition of exceeding safe device pose limits; once triggered, the robot reattempted the task and re-grasped the device in a more favorable joint configuration for further twisting.
Attempting these tasks with the same robot, motion planner, controllers and strategy subtree but without the condition monitor (instead executing a plan made prior to manipulation without adapting to effort observations during execution) resulted in a 0% success rate for all tasks, and effort spikes frequently resulted in catastrophic failure with the Z1 shutting down from overheating or blowing a fuse. This highlights the necessity of adaptive behavior for these tasks.
A limitation of the test setup is that the effort conditions for task success were found empirically from preliminary testing; however, we stress that this is unrelated to the adaptive BT itself, which assumes very little about each condition and would readily accept more sophisticated conditions based on task-space wrench estimation.
IX Conclusion and Future Work
In this work we proposed the adaptive BT, a flexible and highly modular BT design that enables a robot to adapt to an uncertain environment by pairing reactive condition monitoring with adaptive strategy selection. Our test robot used the adaptive BT to safely and reliably complete complex effort-intensive manipulation tasks commonly found in industrial settings, despite using out-the-box controllers. The adaptive BT can be readily added to existing BT implementations without major changes, enhancing robustness while remaining interpretable and without significantly changing the system architecture. Future work will investigate algorithmic and learning-based synthesis of new strategies and conditions for the adaptive BT to enable a robot to generalize to new tasks, including synthesis from more semantic, conceptual descriptions of each strategy, or autonomous discovery of new strategies. We also plan to benchmark against other TAMP methods as well as run experiments beyond the lab.
Acknowledgments
The authors thank Emir Kocak for the implementation and deployment of the motion planner and control interface.
[Actuated Devices in Simulation]
Preliminary testing was done in simulation using Gazebo [49], using simulated devices that are highly accurate to their real-world equivalents, allowing for testing to closely reflect testing on hardware. They can be accessed at [50].

### v3

@@
 This is particularly relevant for effort-sensitive, dynamics-dependent tasks such as tightening a valve.
-To perform these tasks safely and reliably, robots must be able to quickly adapt in response to unexpected changes during task execution.
-Humans can intuitively respond and adapt their manipulation strategy to suit such problems, but representing and implementing such behaviors for robots remains an open question.
-We present the adaptive behavior tree, which enables a robot to quickly adapt to both visual and non-visual observations during task execution, preempting task failure or switching to a different strategy based on data from previous attempts.
-We test our approach on a number of tasks commonly found in industrial settings.
-Our results demonstrate safety, robustness (100% success rate for all but one experiment) and efficiency in task completion (eg, an overall task speedup of 46% on average for valve tightening), and would reduce dependency on human supervision and intervention.
+To perform these tasks safely and reliably, robots must be able to quickly adapt in response to unexpected changes during task execution, and should also learn from past experience to better inform future decisions.
+Humans can intuitively respond and adapt their manipulation strategy to suit such problems, but representing and implementing such behaviors for robots remains a challenge.
+In this work we show how this can be achieved within the framework of behavior trees.
+We present the adaptive behavior tree, a scalable and generalizable behavior tree design that enables a robot to quickly adapt to and learn from both visual and non-visual observations during task execution, preempting task failure or switching to a different manipulation strategy.
+The adaptive behavior tree selects the manipulation strategy that is predicted to optimize task performance, and learns from past experience to improve these predictions for future attempts.
+We test our approach on a variety of tasks commonly found in industry; the adaptive behavior tree demonstrates safety, robustness (100% success rate) and efficiency in task completion (up to 36% task speedup from the baseline).
 I Introduction
 Service robots can revolutionize how entire service-oriented domains work, with potential applications including autonomous maintenance in – often remote – industrial facilities [1, 2], such as oil rigs [3], or care robots in domestic environments [4, 5, 6].
-To complete tasks effectively, robots must use observations from the environment to construct and execute a suitable plan; a number of frameworks have been developed for task and motion planning (TAMP) for manipulation [7, 8, 9, 10].
-This work is motivated by effort-sensitive, dynamics-dependent manipulation tasks for articulated devices commonly found in industrial settings, for example twisting a valve until tightened to a specified torque or pushing an e-stop button until pressed.
+To complete tasks effectively, robots must use observations from the environment to construct and execute a suitable plan; a number of frameworks have been developed for task and motion planning (TAMP) for manipulation [7, 8, 9].
+This work is motivated by effort-sensitive, dynamics-dependent manipulation tasks for articulated devices commonly found in industrial settings, for example twisting a valve until tightened to a specified torque, pushing an e-stop button until pressed, or firmly picking and placing a fragile object.
 Robots generally require highly controlled environments with known task parameters for reliable planning and execution of manipulation tasks; however, the tasks we are interested in pose a significant challenge since they depend on information that can only be determined after sustained physical interaction has already begun.
@@
 For manipulation tasks, different strategies may involve grasping the object differently, or even using a tool to assist manipulation.
-Different strategies can have different trade-offs; for example speed, complexity, effort and power expenditure.
-For example, quickly grasping and twisting a valve handle head-on may be optimal when beginning to twist an initially loose valve, but fail to provide the torque to finish tightening it, while grasping from the side to apply more wrench may be slower for most of the tightening procedure but can fully tighten the valve.
-It may even be that completely different strategies are required to successfully complete different task instances.
-We seek a system to coordinate this strategy selecting and switching behavior.
-We tackle both problems by exploiting the reactivity and modularity of behavior trees.
+Different strategies can have different trade-offs, for example speed, complexity, effort and power expenditure; Fig.
+1 presents examples for tightening a valve.
+We seek a system that can smartly select and switch between known strategies, ideally optimizing for performance by learning from past manipulation attempts.
+Figure 1: Different strategies for tightening a needle valve.
+(1a) Quickly grasping head-on means that all applied torque is generated purely from the wrist joint effort, producing a fast but ‘low-torque’ strategy.
+(1b) Grasping from the side requires more complex motion but enables the robot to use the arm as a wrench, producing a slow but ‘high-torque’ strategy.
+We tackle both problems at the task planning level within the framework of behavior trees, exploiting their reactivity and modularity.
 The key contribution of our work is the adaptive behavior tree, which enables a robot to handle uncertainty in the environment and quickly adapt to unexpected safety-critical changes during task execution, preempting task failure or switching to a more optimal strategy based on data gathered over previous attempts.
-It comprises two stages; a selector that uses data recorded over previous attempts to determine which strategy to deploy, and a monitor that checks a set of conditions concurrently to strategy execution, halting the current strategy if the task is complete or a safety condition is violated and re-attempting the task if necessary.
-The design is highly modular and readily allows for the addition and swapping out of strategy subtrees and condition nodes to suit a new task.
-We test our approach on various case studies representing complex, effort-intensive manipulation tasks for an industrial setting.
+It comprises two stages; a selector that uses data recorded over previous attempts to predict the optimal strategy to deploy, and a monitor that checks a set of conditions concurrently to strategy execution, halting the current strategy if the task is complete or a safety condition is violated and re-attempting the task if necessary.
+The design is highly modular and generalizable, allowing for easy scaling to new strategies and tasks.
+We test the adaptive behavior tree on a diverse range of case studies representing complex, effort-intensive manipulation tasks from industry.
 II Related Work
 II-A Adaptive Control for Robotics
-It may be argued that our problem of manipulating articulated devices with unknown and changing dynamics intrinsics can be solved with an appropriate adaptive controller.
+It may be argued that our problem of manipulating articulated devices with unknown and changing dynamics intrinsics can be solved with an appropriate adaptive controller for each task.
 Adaptive control has existed in the literature for many years, and generally focuses on online parameter estimation and learning for uncertain dynamic systems.
 There exist many good surveys on the literature for adaptive control [10, 11].
-The approach has been extended to robotic manipulation tasks, for example estimating model parameters for an uncertain end-effector payload [13, 14, 15].
+The approach has been extended to robotic manipulation tasks, for example estimating model parameters for an uncertain end-effector payload [12, 13].
 However, we argue that adaptive behavior by parameter tuning is inherently limited in the diversity that it provides for adaptability, and does not readily allow for adaptability to the extent of reactively switching to a completely different manipulation strategy, such as grasping the device in a different way or using a tool, which may be necessary for task completion.
@@
 provide a comprehensive survey of BTs for robotics [14].
-BTs for robotics can be hand-coded [26, 27, 28], planned analytically [29, 30, 31, 32, 33, 34], or even learned from demonstration [35, 36].
+BTs for robotics can be hand-coded [24, 25, 26], planned analytically [27, 28, 29, 30, 31, 32], learned from demonstration [33, 34, 35] or even generated using Large Language Models (LLMs) [36, 37].
 Reactive BTs are particularly relevant to our problem; they use a planning algorithm to automatically create and update a BT on the fly based on observations during task execution, achieving robot behavior that is adaptive to uncertain and changing environments [27, 28].
@@
 When it receives Running, it returns Running and ticks the same child on the next tick, and when it receives Failure, it returns Failure and resets the sequence progress.
-A ReactiveSequence behaves similarly, but when it receives Running, it returns Running and ticks all children up to the current child on the next tick.
-Thus, all prior children are still checked for Failure whenever it is ticked; this allows for reactive behavior.
+A ReactiveSequence is similar, but when it receives Running, it returns Running and also ticks all children up to the current child on the next tick; thus all prior children are still checked for Failure, enabling reactive behavior.
 A Fallback (known as a Selector in other works) propagates the tick to its children from left to right until one of them returns Success, at which point the Fallback stops and returns Success.
@@
 It returns Failure only if all children return Failure.
-A ReactiveFallback behaves similarly, but when it receives Running, it returns Running and ticks all children up to the current child on the next tick.
-Thus, all prior children are still checked for Success whenever it is ticked; this allows for reactive behavior much like in the case of a ReactiveSequence.
+A ReactiveFallback is similar, but when it receives Running, it returns Running and also ticks all children up to the current child on the next tick; thus all prior children are still checked for Success, enabling reactive behavior (much like the ReactiveSequence).
 An ExecutionNode carries out a command, forming the leaves of the BT, and can be of type ActionNode or ConditionNode.
@@
 A ConditionNode checks a condition, returning Success if true or Failure if false.
-A DecoratorNode modifies the return status of a single child node, for example forcing Failure.
-A Precondition is a special type of DecoratorNode that ticks the child only if a particular condition is met, or else returns either Success or Failure (marked by ‘S’ or ‘F’ respectively).
+A DecoratorNode modifies the return status of a single child.
+A Precondition is a special type of DecoratorNode that ticks the child only if a precondition is met, or else returns either Success or Failure (marked by ‘S’ or ‘F’ respectively).
 Figure 2: Categories of BT tree nodes used in this work.
@@
 Fig.
-2 presents the proposed design for an adaptive BT, as well as an example of how it can be integrated within a wider robot system.
-The adaptive BT is composed of two core stages; the strategy selector first decides which strategy to use for the next attempt, based on current perception data as well as data recorded over past attempts (such as joint efforts over time).
+3 presents the adaptive BT, as well as a demonstration of how it can be integrated within a wider robot system.
+The adaptive BT is composed of two core stages; the strategy selector first decides which strategy to use for the next attempt, based on current perception data as well as data recorded over past attempts (such as joint efforts or time taken).
 The condition monitor then wraps around the subtree for the selected strategy, checking a set of task- and strategy-specific conditions concurrently to strategy execution.
 The adaptive BT also admits subtrees to coordinate robot setup prior to strategy selection, and for safely exiting or recovering from the current attempt in preparation for the next.
-Figure 2: Adaptive BT design, integrated within a wider robot system representing the test platform in this work. The coloured lines represent communication between the BT and the rest of the robot.
+Figure 3: The adaptive BT, integrated within a wider robot system representing the test platform in this work.
+The colored lines represent communication lines for perception (blue), motion planning (green) and strategy selection (red).
 The database of the subtrees for all $N\in\mathbb{N}$ known strategies are inserted as modules within a Sequence, but all are preconditioned to only be executed if that specific strategy has been selected.
@@
 The condition checking is such that if the subtree of task-specific conditions returns Success, the task is considered to have been completed successfully, and if it returns Failure, the task is still ongoing.
-This follows the BT design principle of backward chaining, commanding the robot to keep executing the strategy until the task completion conditions are met.
+This follows the BT design principle of backward chaining [15], commanding the robot to keep executing the strategy until task completion conditions are met.
 Meanwhile, if the subtree of strategy-specific conditions returns Failure, a constraint has been violated during execution of the current strategy and the robot must halt the current attempt and try again, potentially selecting a different strategy.
@@
 Each condition is ticked for every tick of the strategy subtree, allowing the robot to rapidly respond to new observations; thus the inherent reactivity of BTs is exploited.
-If a condition is not relevant to the current stage of the strategy subtree, we recommend skipping it by wrapping with an appropriate Precondition.
+If a condition is not relevant to the current stage of the strategy subtree, it can be skipped by wrapping with an appropriate Precondition.
 The subtrees for exit and recovery allow the robot to safely exit the current attempt, due to task success or preempting failure respectively.
@@
 The only required knowledge for adding the proposed design into a larger BT is that the adaptive BT must be repeatedly ticked while Running, and will return Success if the task was completed successfully and Failure otherwise.
-To avoid getting stuck, we recommend limiting the adaptive BT to a finite number of allowed attempts.
+The adaptive BT is designed to coordinate efficient selection, monitoring and switching of strategies, not the synthesis of the strategies themselves.
+The adaptive BT would synergize well with methods for the automatic synthesis of BTs [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] to generate new strategies for it to use, but exploring this is beyond the scope of this work.
 V Adaptive BT for Manipulation
-We now apply the proposed adaptive BT to the context of manipulation tasks wherein the robot must manipulate an articulated device; task success occurs when the device is considered sufficiently manipulated.
-For all manipulation tasks explored in this work, the manipulation procedure consists of identifying the pose of the device, then approaching, grasping, manipulating and finally retracting from the device.
+We now apply the adaptive BT to manipulation tasks; Fig.
+4 presents the general subtrees for all strategies used in this work.
+We stress that these designs are just examples used to test the adaptive BT, with an emphasis on simplicity; the adaptive BT readily accepts new strategies for new tasks.
+Figure 4: Subtrees for strategies used in this work. (4a) Manipulating an articulated device. (4b) Grasping an object. (4c) Moving an object to a goal.
+Manipulating Articulated Devices.
 Fig.
-3 presents the designs used for each subtree of the adaptive BT.
-We stress that these designs are just examples used to test the adaptive BT, with an emphasis on simplicity.
-Figure 3: Subtrees used for all manipulation tasks considered in this work.
-Setup.
-The robot moves to a pose suitable for observing the task instance to determine which strategy to select, and then estimates the pose of the device handle.
-If the device handle has rotational symmetry and thus has more than one valid pose, the robot later uses the most suitable pose that is within safe device handle pose limits for the selected strategy.
-Strategy Subtree.
-The robot approaches, grasps and manipulates the device with its end-effector.
-These motions are each defined by a waypointed end-effector pose trajectory that is solved using the robot’s motion planner, with the planning solution then sent to the robot’s controllers.
-The pose trajectories are different between strategies, enabling the robot to manipulate the device in different ways.
-While the strategies tested in this work all use the same subtree design, this is not a limitation of the adaptive BT, and in general each strategy can use a completely different design, enabling greater variety between strategies.
-Task-Specific Conditions.
-If currently manipulating the device, the robot checks whether the device is considered sufficiently manipulated.
-This condition is specified differently depending on the particular task; for example, when tightening a valve, the current joint efforts are used to estimate tightness, while when pressing a button, the joint efforts are used to estimate whether it has been pressed.
-Strategy-Specific Conditions.
-If approaching, grasping or manipulating the device, the robot checks that its end-effector pose is within tolerance of the current pose trajectory.
-When manipulating the device, the robot checks that its joint efforts do not exceed the safe limits for the strategy.
-The robot also checks that the device handle pose is still within safe limits for continuous manipulation using the strategy.
-This is important since a task like tightening a valve may require multiple revolutions of the handle, yet the robot may not be physically able to do this in one motion due it its own joint limits.
-Thus, this condition tells the robot when it should stop and re-grasp the device handle, preventing the robot from overextending itself during manipulation.
-Recovery/Exit.
-Both subtrees use the same design; the robot retracts its end-effector from the device (if grasped) and then resets to its starting pose and joint configuration.
-The retraction is also defined by a waypointed end-effector pose trajectory and solved via motion planning.
+4a presents general subtree structures for strategies that manipulate an articulated device, where task success occurs when the device is considered sufficiently manipulated (e.g.
+valve is tight).
+The strategies consist of approaching, grasping and manipulating the device (e.g.
+twist by one-half rotation), while checking safety conditions.
+Note that these same subtree structures are used for both the low-torque and high-torque strategies depicted in Fig.
+1, and even the strategies for manipulating the different devices (e.g.
+tightening a valve vs pressing a button), though the action and condition nodes have different parameters (or even implementations) for each strategy.
+Object Grasping.
+Fig.
+4b presents general subtree structures for strategies that grasp an object, where task success occurs when the object is considered firmly grasped (e.g.
+by checking gripper effort).
+The strategies consist of approaching and grasping, while checking safety conditions.
+Object Placing.
+Fig.
+4c presents general subtree structures for strategies that move an (already grasped) object to a goal pose, where task success occurs if the strategy completes without preemption (thus no task-specific conditions).
+The strategies consist of moving and then releasing the object, while checking safety conditions.
+Setup/Recovery/Exit. In this work, these are simply action nodes that position the robot to view the scene for strategy selection (for setup) or reset the robot pose (otherwise).
+Strategy Selection.
+The SelectStrategy node estimates the optimal manipulation strategy to use; for this work, the optimal strategy $a^{*}$ for task state $s$ minimizes task duration $t_{\mathrm{est}}$ while not violating strategy-specific constraints (often corresponding to the safety conditions, e.g.
+maximum safe joint efforts).
+This is equivalent to the following problem:
+$a^{*}(s)=\min_{a\in\mathcal{A}_{\mathrm{task}}}t_{\mathrm{est}}(a,s)\quad\text{subject to}\quad a\in\mathcal{A}_{\mathrm{safe}}(s),$
+where strategies $a$ are selected from the set of available task-relevant strategies $\mathcal{A}_{\mathrm{task}}$, and $\mathcal{A}_{\mathrm{safe}}\subseteq\mathcal{A}_{\mathrm{task}}$ is the set of feasible strategies given the constraints.
+Both $t_{\mathrm{est}}$ and $\mathcal{A}_{\mathrm{safe}}$ are estimated from data recorded from previous attempts; if no data exists at $s$, we let $\mathcal{A}_{\mathrm{safe}}=\mathcal{A}_{\mathrm{task}}$, and calculate $t_{\mathrm{est}}$ from nominal durations of the strategy’s motion trajectories.
 VI Implementation
 Test Platform.
-We used an ANYbotics ANYmal C quadruped[38] with mounted Unitree Z1 robotic arm[39], equipped with a wrist-mounted Intel Realsense Depth Camera D345[40] for perception.
+We used an ANYbotics ANYmal quadruped with mounted Unitree Z1 robotic arm, equipped with a wrist-mounted Intel Realsense Depth Camera D345 for perception.
 ROS, the Robot Operating System, was used for the implementation of the test setup as well as communication with hardware [39].
 Behavior Tree.
-BT.CPP 4.0 was used to construct the adaptive BT.
+BT.CPP 4.0 [38] was used to construct the adaptive BT.
 The ActionNode types specific to this work take the form of ROS action and service clients, implemented as derived classes of the StatefulActionNode from the BT.CPP core library; this node type supports asynchronous actions and is recommended for this use case.
-Strategy Selection.
-The SelectStrategy node estimates the optimal manipulation strategy.
-For this work, the optimal strategy $a^{*}$ completes the task in the shortest time while being feasible given the effort data recorded during manipulation.
-This is equivalent to the following problem:
-$a^{*}=\min_{a\in\mathcal{A}_{\mathrm{tsk}}}t_{\mathrm{est}}(a)\quad\text{s.t.}\quad T_{\mathrm{max},i}(a)<T_{\mathrm{lim},i}(a)\;\forall i$ (1)
-where strategies $a$ are selected from the set of available task strategies $\mathcal{A}_{\mathrm{tsk}}$, $t_{\mathrm{est}}(a)$ is the estimated time for task completion, $T_{\mathrm{max},i}(a)$ are the largest recorded absolute effort values for each arm joint $i$, recorded from previous manipulation attempts using $a$, and $T_{\mathrm{lim},i}(a)$ are the maximum permitted absolute effort values for each joint $i$ under strategy $a$.
-The effort data used for strategy selection is recorded from the arm joint efforts during each manipulation attempt, and stored in a database accessible to SelectStrategy.
 Motion Planning.
@@
 The planner solves an EXOTica DynamicTimeIndexedShootingProblem, tracking pre-defined waypointed end-effector pose trajectories for the approach, grasp, manipulate and retract stages of the manipulation task for each strategy.
-The pose trajectories are defined relative to the pose of the device handle; in this way, the planner is generalizable to arbitrary device poses and handle angles (so long as robot joint limits permit the motion).
+The pose trajectories are defined relative to the pose of the object or device handle; in this way, the planner generalizes to arbitrary poses and handle angles (so long as robot joint limits permit the motion).
 In the case of planning failure, the planner action nodes return Failure, prompting the adaptive BT to retry the task.
 Control.
-This work uses the standard controllers provided with the ANYmal C and Unitree Z1.
-These controllers are by no means perfect, and in practice there was often non-negligible deviation between the true end-effector pose and that in the motion plan, especially when the arm was outstretched; to ensure safe operation, this deviation is checked by the condition monitor during manipulation.
-This is a major cause of the re-attempts that would occasionally occur during testing as seen in section VIII.
-Using a bespoke whole-body controller that compensates for both the motion of the base and arm would improve performance, but this is beyond the scope of this work; instead, we demonstrate how our adaptive BT can be used to overcome these problems by autonomously preempting failure and retrying.
+This work uses the standard controllers provided with the ANYmal and Unitree Z1.
+The combined controller is by no means perfect, since the ANYmal controller has no awareness of the arm and cannot compensate for the induced wrench on the quadruped when the arm is outstretched.
+In practice, there was often non-negligible deviation between the true end-effector pose and that in the motion plan, especially when the arm was outstretched; to ensure safe operation, this deviation is checked by the condition monitor during manipulation.
+Using a bespoke whole-body controller that compensates for both the motion of the quadruped and arm would improve performance, but this is beyond the scope of this work; instead, we demonstrate how our adaptive BT can be used to overcome such problems at the task planning level by autonomously preempting failure and retrying.
 Perception.
-An Apriltag [43] was used to identify the pose of the device handle in an ‘unrotated’ state, and invariant template matching [44] was used to identify the handle angle (and thus the pose of the rotated handle).
-VII Experiments
+Apriltags [41] were used to identify the pose of objects as well as device handles in an ‘unrotated’ state, and invariant template matching [42] was used to identify handle angles and the pose of the rotated handles.
+VII Case Studies
+All device manipulation tasks use the general subtree structures depicted by Fig.
+4a, albeit with task- and strategy-specific implementations of the action and condition nodes.
+Appendix A describes estimation of $s$ and $\mathcal{A}_{\mathrm{safe}}$ for all tasks.
 Needle Valve.
-The robot twists a Tameson NLS-012 needle valve[45] from unknown starting tightness until tight.
-Two strategies were provided; the faster, simpler ‘low-torque’ strategy grasps the handle head-on while the slower, more complex ‘high-torque’ strategy grasps it from the side.
-Fig.
-4 depicts these strategies.
-The experiment is conducted using each strategy on its own, as well as using both strategies such that the BT can adaptively select between them.
-To demonstrate the modularity offered by the adaptive BT, we also consider a walk-and-manipulate task where the robot must navigate to, tighten, and navigate away from the needle valve.
-The adaptive BT is inserted as a subtree within a larger BT that also commands navigation to and from the device (using the standard ANYmal C navigation and locomotion software).
-Fig.
-5a-5c illustrates this experiment.
-Figure 4: An annotated timeline of tightening the needle valve using the adaptive BT with both available strategies, with plots of the effort in the wrist joint during each manipulation attempt A, B, C and D.
+The robot twists a Tameson NLS-012 needle valve from unknown starting tightness until tight, and loosens it to a user-specified angle, with access to the two strategies depicted by Fig.
+1.
+To demonstrate how the adaptive BT can be readily placed within larger BT structures, we also consider a walk-and-manipulate task where the robot navigates to, tightens, and navigates away from the needle valve (with access to both manipulation strategies, but the required applied torque for tightness reduced so as to be achievable by the ‘low-torque’ strategy).
+The adaptive BT is inserted as a subtree within a larger BT that also commands navigation to and from the device (using the standard ANYmal navigation and locomotion software).
+Globe Valve. The robot twists a Dikkan DN40 PN16 globe valve from unknown starting tightness until tight, with access to a single strategy depicted by Fig. 5b.
+E-Stop Button.
+The robot manipulates an IDEC HW1X-BV411-R emergency stop; it pushes the button until fully pressed, with access to a single strategy depicted by Fig.
+5c, and twists it until released, with access to a single strategy very similar to that depicted by Fig.
+1a.
+Groceries Packing.
+The robot packs a plastic bottle, (full) juice box, (real) orange and (plastic) banana into a basket, in a user-specified order, as depicted by Fig.
+5d.
+The grasping stage uses a single strategy, with subtree structures depicted by Fig.
+4b, determining firm grasping by checking gripper joint effort.
+The packing stage has access to a continuous set of strategies; all strategies have the same subtree structures depicted by Fig.
+4c, but the speed of the motion trajectories is parameterized by a speed factor (i.e.
+a strategy) $a\in\mathcal{A}_{\mathrm{safe}}$, where $\mathcal{A}_{\mathrm{safe}}$ is determined by the weight of the object (lower maximum safe speeds for heavier objects, which may slip out of grasp when the motions are too fast).
+The adaptive BTs for the two stages are placed in sequence such that the robot returns to the first stage if the second stage is preempted.
+Figure 5: Snapshots from experiments.
+(5a) Walking to the needle valve (as part of the walk-and-manipulate task).
+(5b) Tightening the globe valve.
+(5c) Pushing the e-stop button until fully pressed.
+(5d) Groceries packing.
+Accompanying video with experiments: https://youtu.be/im7OCkEiQsU.
+Baseline.
+For all tasks, we use backward chaining with stacked Fallbacks as the baseline BT structure [15].
+This is the standard BT design used when the robot must choose between different actions with robustness to failure.
+For fair comparison to the adaptive BT, we manually arranged the (nominal) ‘faster’ strategies before ‘slower’ ones in the baseline’s Fallbacks such that the robot tries them first, with a reminder that the adaptive BT automates (and optimizes) the strategy selection process.
+Since the baseline cannot select from a continuous set of strategies, it can only (safely) use the lowest speed factor when grocery packing.
+VIII Results and Discussion
+To evaluate the adaptive BT, we performed a series of ablations to assess how each component impacts performance, with comparison to the baseline where relevant.
+For our discussion, we refer to the results in Table I, which records, for 10 trials of each experiment, the cumulative success rate after each attempt (ignoring re-attempts triggered to change strategy and re-grasping when device handle pose limits were exceeded) and time taken for task completion.
+A maximum of 5 attempts was permitted for each trial, task success was verified by checking the scene after the robot had exited the task, and the initial device states were reset to be the same for each trial.
+Strategy effort limits were set to the Z1’s maximum safe operating efforts scaled down by a safety factor, except for the needle valve low-torque strategy, which used a larger safety factor to avoid the quadruped becoming unstable from induced wrench during heavy twisting.
+TABLE I: Experiment Results for All Tasks
+Task Experiment Cumulative Success Rate (%) After Attempt Number Time Taken (s)
+1 2 3 4 5 Minimum Mean
+Ablation A: RFP Enabled (SAT + LPE Disabled, i.e. Equivalent to Baseline)
+Needle Valve (Tighten) Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 135 (Fail) 136 (Fail)
+High-Torque Strategy Only 50 70 90 100 Done 279 336
+Walk-and-Manipulate 60 80 90 100 Done 212 230
+Needle Valve (Loosen) Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 58 (Fail) 58 (Fail)
+High-Torque Strategy Only 70 90 100 Done Done 228 247
+Globe Valve Tighten 70 100 Done Done Done 95 107
+E-Stop Button Press 80 100 Done Done Done 28 32
+Release 70 80 100 Done Done 45 63
+Groceries Packing Lowest Speed Only 60 70 90 100 Done 261 269
+Ablation B: RFP + SAT Enabled (LPE Disabled)
+Needle Valve (Tighten) Both Strategies (Baseline) 70 100 Done Done Done 175 186
+Both Strategies (Adaptive BT) 80 100 Done Done Done 175 182
+Needle Valve (Loosen) Both Strategies (Baseline) 60 90 100 Done Done 259 283
+Both Strategies (Adaptive BT) 80 100 Done Done Done 204 212
+Groceries Packing Lowest Speed Only (Baseline) 60 70 90 100 Done 261 269
+Weight-Based Speed (Adaptive BT) 50 70 80 100 Done 196 206
+Ablation C: RFP + SAT + LPE Enabled (Complete Adaptive BT)
+Needle Valve (Tighten) Both Strategies (Baseline) 80 100 Done Done Done 94 126
+Both Strategies (Adaptive BT) 90 90 100 Done Done 58 118
+Needle Valve (Loosen) Both Strategies (Baseline) 60 90 100 Done Done 259 283
+Both Strategies (Adaptive BT) 80 100 Done Done Done 172 182
+Reactive Failure Preemption (RFP).
+For Ablation A, the robot only has access to one strategy (preventing strategy adaptation) and manipulation data is reset between trials (preventing learning from past experience).
+We see a 100% success rate within the permitted number of attempts for all but one experiment when using the adaptive BT, despite having no knowledge of the valves’ initial tightness, or how much they needed to be twisted to be tightened.
+The robot never suffered catastrophic failure since the condition monitor preempted failure before it occurred.
+When using the low-torque strategy, the robot aborted the needle valve tasks after exceeding the safe effort limit for the arm’s wrist joint.
+This demonstrates the safety enabled by the adaptive BT, since the robot was able to react and safely exit the task before overexerting itself.
+The robot was also able to safely twist the valves for an arbitrary number of rotations without the danger of overextending itself, by periodically re-grasping the valves using the condition of exceeding safe device pose limits; once triggered, the robot re-attempted the task and re-grasped the device in a more favorable joint configuration for further twisting.
+The robot was also able to firmly grasp the different-shaped fragile objects without damaging them by reacting to spikes in gripper effort.
+Under this ablation, the adaptive BT produces equivalent behaviors to the baseline.
+However, the adaptive BT enables easier scalability of task composition over the baseline, since the adaptive BT provides an abstracted, modular interface for appending and removing subtrees for strategies and conditions to the strategy database (see Fig.
+3), while editing the baseline backward chained BT structure can become unwieldy as the number of strategies and conditions grows.
+Strategy Adaptation (SAT).
+For Ablation B, the robot has access to all strategies but manipulation data is still reset between trials (preventing learning from past experience).
+We see substantial average task speedups for needle valve tightening (46%) and loosening (14%) as well as groceries packing (23%) when using the adaptive BT compared to Ablation A; the cause for this speedup is that the adaptive BT selects faster strategies when feasible and switches to slower, ‘safer’ strategies when necessary.
+For needle valve tightening, since the low-torque strategy is approximately 5 times faster while twisting, it is optimal for the robot to use it for most of the manipulation sequence, but when the robot experiences a sudden effort spike as the valve begins to tighten, it quickly adapts and uses the high-torque strategy to fully tighten the valve; Fig.
+6 provides an annotated timeline.
+We note that this is similar to behavior displayed by humans, twisting quickly when the valve is loose but changing grip near the end to be able to apply more wrench without hurting oneself.
+For this particular task the baseline produces the same behavior, leading to roughly the same average time taken; this is because the optimal behavior for this task (start low-torque, switch to high-torque at the end) can also be achieved using a simple Fallback.
+However in general this is not the case, as demonstrated by the needle valve loosening task, wherein the adaptive BT achieves a 25% average speedup over the baseline.
+For this task, both the adaptive BT and baseline first (unsuccessfully) try the low-torque strategy (which cannot safely apply enough torque to loosen the valve) and then switch to the high-torque strategy for the first successful twist, after which the valve is loose enough to be loosened further using the low-torque strategy.
+While the logic of the Fallback means that the baseline will keep using the high-torque strategy for the remainder of the task, the adaptive BT optimizes over feasible strategies at each re-attempt, and thus chooses the (faster) low-torque strategy for the remainder of the task.
+Figure 6: An annotated timeline of tightening the needle valve using the adaptive BT with both available strategies, with plots of the effort in the wrist joint during each manipulation attempt A, B, C (which all use the ‘low-torque’ strategy) and D (which uses the ‘high-torque’ strategy).
+The adaptive BT initially selects the low-torque strategy to twist quickly while the valve is loose but switches to the high-torque strategy near the end to apply more wrench and finish tightening.
 Note the effort spikes during attempts C and D; the first indicates that the low-torque strategy is no longer feasible, while the second indicates that the device is sufficiently tightened.
 For clarity, only the lowest active effort limit is shown in each plot.
-Figure 5: Snapshots of experiments performed.
-(5a) Navigating to the needle valve.
-(5b) Tightening the needle valve.
-(5c) Navigating away from the needle valve.
-(5d) Pushing the e-stop button until fully pressed.
-(5e) Twisting the e-stop button until released.
-(5f) Tightening the globe valve.
-E-Stop Button.
-The robot manipulates an IDEC HW1X-BV411-R emergency stop[46]; it pushes the button until fully pressed and twists it until released.
-One strategy was provided for each task; Fig.
-5d and 5e depict these strategies.
-A couple thin foam strips were placed on the Unitree Z1’s gripper to allow it to tightly grasp the button for twisting.
-Globe Valve.
-The robot twists a Dikkan DN40 PN16 globe valve[47] from unknown starting tightness until tight.
-One strategy was provided for this task; Fig.
-5f depicts this strategy.
-The valve’s diaphragm and O-ring were removed to make it loose enough to be twisted by the Unitree Z1, and some thin foam strips were placed on the handle to allow the Z1’s gripper to tightly grasp the handle.
-VIII Results and Discussion
-An accompanying video for the experiments performed is provided with this work [48].
-Each experiment was repeated for 10 trials with the device handle pose and recorded effort data reset for each trial.
-Strategy effort limits were set to the Unitree Z1’s maximum safe operating efforts scaled down by a small safety factor, except for the low-torque strategy which used a larger safety factor to avoid the base becoming unstable from the arm’s wrench during heavy twisting.
-Task success was verified by checking the device after the robot had exited the task.
-Table I records, for each experiment, the cumulative success rate after each attempt, ignoring re-attempts triggered to change strategy and re-grasping when device handle pose limits were exceeded, as well as times taken for task completion.
-A maximum of 5 attempts was permitted for each trial.
-Fig.
-4 provides an annotated timeline of tightening the needle valve when having access to both associated strategies.
-TABLE I: Results for All Manipulation Tasks Using the Adaptive BT
-Device Experiment Cumulative Success Rate (%) After Attempt Number Time Taken (s)
-1 2 3 4 5 Minimum Mean
-Needle Valve Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 135 (Fail) 136 (Fail)
-High-Torque Strategy Only 50 70 90 100 Done 279 336
-Both Strategies (Adaptive) 80 100 Done Done Done 175 182
-Walk-and-Manipulate 60 80 90 100 Done 212 230
-E-Stop Button Press 80 100 Done Done Done 28 32
-Release 70 80 100 Done Done 45 63
-Globe Valve Tighten 70 100 Done Done Done 95 107
-Table I shows that the robot achieved a 100% success rate within the permitted number of attempts for all but one experiment when using the adaptive BT, despite having no knowledge of the valves’ initial tightness, or how much they needed to be twisted to be tightened.
-The robot never suffered catastrophic failure since the condition monitor preempted failure before it occurred.
-When only having access to the low-torque strategy, the robot aborted tightening the needle valve after the measured joint effort exceeded the strategy effort limit for the arm’s wrist joint.
-This demonstrates the safety enabled by the adaptive BT, since the robot was able to react and safely exit the task before overexerting itself, and would not try again because the manipulation data suggested that no feasible strategies were available.
-However, the utility of having access to the low-torque strategy is made clear when comparing the task completion times for using the high-torque strategy only and adaptively using both strategies.
-Since the low-torque strategy is approximately 5 times faster while twisting, it is optimal for the robot to use it for most of the manipulation sequence, but when the robot experiences a sudden effort spike as the valve begins to tighten, it quickly adapts and uses the high-torque strategy to fully tighten the valve; this resulted in an overall task speedup of 46% on average.
-We note that this is similar to behaviour displayed by humans, twisting quickly when the valve is loose but changing grip near the end to be able to apply more wrench without hurting oneself.
-The robot was also able to safely twist the valves for an arbitrary number of rotations without the danger of overextending itself.
-This was done by periodically re-grasping the valves using the condition of exceeding safe device pose limits; once triggered, the robot reattempted the task and re-grasped the device in a more favorable joint configuration for further twisting.
-Attempting these tasks with the same robot, motion planner, controllers and strategy subtree but without the condition monitor (instead executing a plan made prior to manipulation without adapting to effort observations during execution) resulted in a 0% success rate for all tasks, and effort spikes frequently resulted in catastrophic failure with the Z1 shutting down from overheating or blowing a fuse.
-This highlights the necessity of adaptive behavior for these tasks.
+Learning from Past Experience (LPE).
+For Ablation C, the robot has access to all strategies as well as the manipulation data recorded from Ablations A and B; note that the baseline has no mechanism for learning from past manipulation data.
+For needle valve loosening, the adaptive BT achieves an even greater average speedup (36%) over the baseline; this is because past manipulation data suggests that the valve will initially be too tight to be loosened with the low-torque strategy, so the robot has learned to start with the high-torque strategy (while in Ablation B, due to lack of knowledge of the valve’s initial tightness, it would initially try low-torque then immediately switch to high-torque).
+For needle valve tightening, we see no change from Ablation B if the robot has no initial estimate of the required rotation until tight, i.e.
+task state $s$.
+Note that as we vary initial $s$, in some cases it is faster to start with the high-torque strategy, since it avoids the overhead of switching from low-torque to high-torque to finish tightening.
+If we now provide the adaptive BT with an accurate initial estimate for $s$, it always starts with the strategy that minimizes time taken to tighten the valve according to past experience (unlike Ablation B, which always starts with low-torque).
+This is illustrated by Fig.
+7, which plots minimum time taken to tighten the needle valve for 5 different known required rotations, with the adaptive BT achieving up to 34% speedup over the baseline (at 0.5 radians).
+For Ablation C’s results for this task in Table I, the 10 trials are instead 2 trials each over these 5 different known required rotations; here the adaptive BT achieves 6.3% average speedup over the baseline.
+In practice, initial $s$ could be a random variable with a known distribution; for this case, the mean time taken over this distribution would be the appropriate $t_{\mathrm{est}}$ to use for strategy selection.
+If the distribution is unknown, it could be estimated empirically over many manipulations.
+Figure 7: Minimum time to tighten the needle valve for different (known) required rotations until tight.
+The adaptive BT starts with the strategy that minimizes time taken to tighten the valve according to past experience.
 A limitation of the test setup is that the effort conditions for task success were found empirically from preliminary testing; however, we stress that this is unrelated to the adaptive BT itself, which assumes very little about each condition and would readily accept more sophisticated conditions based on task-space wrench estimation.
 IX Conclusion and Future Work
-In this work we proposed the adaptive BT, a flexible and highly modular BT design that enables a robot to adapt to an uncertain environment by pairing reactive condition monitoring with adaptive strategy selection.
-Our test robot used the adaptive BT to safely and reliably complete complex effort-intensive manipulation tasks commonly found in industrial settings, despite using out-the-box controllers.
-The adaptive BT can be readily added to existing BT implementations without major changes, enhancing robustness while remaining interpretable and without significantly changing the system architecture.
+In this work we proposed the adaptive BT, a flexible and highly modular BT design that enables a robot to adapt to environment uncertainties by pairing reactive condition monitoring with adaptive strategy selection, and learning from past experience to predict the optimal strategy given the task state.
+Our test robot used the adaptive BT to safely and robustly complete a diverse range of complex effort-intensive manipulation tasks commonly found in industrial and domestic settings, despite using out-the-box controllers, and with reduced time taken over the baseline BT design for all experiments where multiple strategies were available.
+The adaptive BT can be readily added to existing BT implementations without major changes, enhancing performance while remaining interpretable and without significantly changing the system architecture.
 Future work will investigate algorithmic and learning-based synthesis of new strategies and conditions for the adaptive BT to enable a robot to generalize to new tasks, including synthesis from more semantic, conceptual descriptions of each strategy, or autonomous discovery of new strategies.
-We also plan to benchmark against other TAMP methods as well as run experiments beyond the lab.
-Acknowledgments
-The authors thank Emir Kocak for the implementation and deployment of the motion planner and control interface.
-[Actuated Devices in Simulation]
-Preliminary testing was done in simulation using Gazebo [49], using simulated devices that are highly accurate to their real-world equivalents, allowing for testing to closely reflect testing on hardware.
-They can be accessed at [50].
+Appendix A Computing The Feasible Set of Strategies
+When manipulating articulated devices, all feasible strategies for task state $s$ are such that the robot arm joint efforts are not exceeded when using those strategies in the window $\{s^{\prime}:s_{L}\leq s^{\prime}-s\leq s_{U}\}$ around $s$.
+Thus, $\mathcal{A}_{\mathrm{safe}}$ is given by
+[…]
+$\forall i\in[1..N_{\mathrm{DOF}}],s_{L}\leq s^{\prime}-s\leq s_{U}\},$
+where $T_{\mathrm{est},i}(a,s)$ are the estimated efforts for each arm joint $i$ at task state $s$ using strategy $a$, recorded from previous manipulation attempts, and $T_{\mathrm{lim},i}(a)$ are the maximum safe efforts for each joint under strategy $a$.
+$T_{\mathrm{est},i}$ are recorded from the joint efforts during each manipulation attempt.
+For valve manipulation, the task state $s$ is the total rotation angle $\theta$ away from being tight, with initial estimate 0 radians for loosening tasks.
+For tightening tasks, initial $\theta$ is estimated to be a large default value if unknown, but once the valve is tight, all $\theta$ values in the new batch of manipulation data are shifted such that the final entry is 0 radians.
+For button manipulation, the task has a (known) binary on/off state.
+When grocery packing, the speed factor is computed from the additional effort that the weight of the object induces in the arm joints. Accordingly, $\mathcal{A}_{\mathrm{safe}}$ is given by
+$\mathcal{A}_{\mathrm{safe}}(s)=\{a\in\mathbb{R}^{+}:a_{L}\leq a\leq a_{H}(s)\}$
+where $\text{where }a_{H}(s)=\min\bigl{\{}a_{L},1-\sum_{i}\alpha_{i}|T_{\mathrm{obj},i}(s)-T_{\mathrm{pre},i}|\bigr{\}},$
+where $T_{\mathrm{pre},i}$ and $T_{\mathrm{obj},i}(s)$ are the efforts recorded immediately before and after lifting the object respectively, $a_{L}$ is a minimum speed factor and $\alpha_{i}\geq 0$ are pre-tuned parameters.
+State $s$ is the object weight, but is not directly measured.